Harriet's age is 3 more than 2 times Angela age. If the sum of their ages is 51 how old is each girl?
Answer by Boreal(15235)

Angela's age=x
Harriet's age=2x+3
They add to 51
x+2x+3=51
3x+3=51
3x=48
x=16 Angela
2x+3=35 Harriet
They add to 51.